OpenStudy (anonymous):

The diagonals of a square are 4 meters long. The side of this square is equal to the diagonal of a second square. Find the side of the second square.

OpenStudy (mathteacher1729):

1. Have you tried drawing a sketch? 2. Do you know how to use the Pythagorean Theorem?

OpenStudy (anonymous):

Yes, but I don't really know how to draw the picture.

OpenStudy (mathteacher1729):

Do you know how to draw a square? :)

OpenStudy (anonymous):

I don't get what it means by the of this square is equal to the diagonal of the second square.

OpenStudy (mathteacher1729):

It means the length of the side of one square is equal to the length of the diagonal of another square. It's comparing lengths. Without a picture, it can be very confusing.

OpenStudy (mathteacher1729):



OpenStudy (mathteacher1729):

How did you get 8?

OpenStudy (mathteacher1729):

How did you get 2.8?

OpenStudy (anonymous):

2

OpenStudy (mathteacher1729):

So I'm asking my questions to try to understand your thought process. :) If you just write numbers... I have no way of knowing how you're getting those numbers, and I can't be much help.

OpenStudy (mathteacher1729):

This problem doesn't actually require any computations at all. (though I'm a little concerned that you don't know pythag theorem ... it appears everywhere).

OpenStudy (anonymous):

Using the Pythagorean Theroem a^2+b^2=c^2 a^2+b^2=8 a=b=2

OpenStudy (mathteacher1729):